A block of mass m_1 = 1.5 kg is placed on a horizo
Answer #1

equate the forces

m2g - T = (m1+m2)a ----------

3.5* 9.81 - T = 4*5.2

T = 13.535

now,

T - mu*m1g = m1*a ---------1

(13.535 - 1.5*5.2)5.2*9.8 = mu

mu = 0.112